OpenStudy (anonymous):

An electron travels 1.64 m in 4.32 × 10^−8 s. How fast is it traveling? Answer in units of m/s

OpenStudy (anonymous):

=1.64/(4.32*10^-8) =1.64*10^8/4.32 =3.7962*10^7 m/s

OpenStudy (anonymous):

Thanks!